# MESA: Matching Everything by Segmenting Anything

### 30 Jan 2024 v1

@@ 1 Introduction
The underlying goal of feature matching is to find correspondences of each point in the source image from the target image. When considering a specific source point, the target image is initially regarded as a candidate correspondence set, encompassing all the points, which is highly redundant given that the default goal is to find a single point match. Therefore, feature matching essentially involves reducing the matching redundancy, which is caused by troublesome matching noises. The classical sparse matching framework [29, 13] reduces the matching redundancy from entire image to keypoints by feature detection. However, achieving reliable keypoints has been challenging [28, 18], resulting in a decline in performance. To avoid the detection failure, semi-dense methods [8, 38], also known as the detector-free methods, are proposed. These methods search for matches densely at the coarse level and refine precise matches locally in the fine level. Although precision improvements are obtained, they rely on heavy feature comparison for matching redundancy reduction. Hence, limited input resolution and error-prone feature computation in matching redundancy decrease their accuracy. Meanwhile, dense methods [16, 25], which produce a warp with confidence for the input image pair, also suffer from precision degradation associate with redundant global computation. Thus, superior matching redundancy reduction is urgently required for both semi-dense and dense matching methods.
@@
On the other hand, most matching redundancy can be effectively and efficiently identified through high-level image understanding and only strongly correlated local areas (also known as regions) need dense feature comparison to determine precise matches. (Fig. 1). Therefore, some methods [14, 9] perform learning-based overlap segmentation before matching. However, implicit learning needs non-reusable computational overhead and the redundancy still exists in the overlap. To address these issues, some work turn to explicit semantic prior [17, 49]. Unlike manually specifying a topic number to group feature patches [17], Area to Point Matching (A2PM) framework proposed by SGAM [49] provides a more intuitive way to reduce matching redundancy. Specifically, this framework establishes semantic area matches first where matching redundancy is largely removed, and then obtains accurate correspondences within these areas utilizing an off-the-shelf point matcher. Nevertheless, SGAM relies on semantic segmentation to determine area matches. Its performance, thus, decreases when encountering inexact semantic labeling and semantic ambiguity [49]. Furthermore, the benefits of A2PM cannot be extrapolated to more general scenes due to the close-set semantic labels. Hence, reducing matching redundancy through semantic segmentation suffers from impracticality.
@@
In this work, we propose Matching Everything by Segmenting Anything (MESA, Fig. 2), a method for precise and robust area matching from the SAM segmentation. It mainly contains a novel multi-relational graph and a graphical area matching technique. Since independent areas only contain local information, it is critical to model the global context for exact area matching. Thus we propose a novel multi-relational graph [1] structure, named Area Graph (AG), which initially takes SAM areas as nodes and connects them based on two different inter-area spatial relations. To accommodate scale variation and enable more efficient matching, we generate more nodes at various image scales and incorporate corresponding edges by graph completion. Then we probabilize the area matching and derive two distinct graphical models from the AG: Area Markov Random Field (AMRF) and Area Bayesian Network (ABN). Subsequently, the area matching is formulated as an energy minimization problem on the AMRF. By carefully designing the energy function, we make this energy minimization effectively solvable through the Graph Cut algorithm [22]. To calculate the graph energy, we introduce a learning-based model to calculate area similarities by patch-level classification, enhancing the area matching interpretability. The conditional independence in ABN is also leveraged to facilitate energy calculation. Finally, to fully consider global structures in both AGs of the input image pair, we propose a global matching energy. After minimizing this energy, we achieve accurate and robust area matching, which allows for fine-grained point matching within area matches, improving the matching performance.
@@
Figure 2: Overview of MESA. We first construct ❷ Area Graphs from ❶ SAM segmentation. Then the graph is turned to two graphical models based on its two different edges. The area matching is formulated as an energy minimization through ❸ Area Markov Random Field, which is effectively solvable by ❼ Graph Cut. The prerequisite ❻ Graph Energy Calculation is efficiently realized by ❺ Area Bayesian Network and our ❺ Learning Area Similarity Calculation. Finally, ❽ Global Energy Minimization determines the best area match, which serves as the input of subsequent point matcher for precise feature matching, following the ❾ Area to Point Matching framework [49].
@@ 3 Area Graph
SAM [21] has the capability to accurately delineate multiple areas within the image. These areas possess implicit semantic due to the high-level image understanding in SAM, hereby favoring precise point matching within them [49]. However, direct area matching on SAM results is unacceptable, because global information is ignored in independent areas and fixed area sizes hinder robust matching under scale changes. Therefore we propose using a novel graph structure, named Area Graph (AG), to capture the structural information of image areas and construct their scale hierarchy, contributing to accurate and robust area matching.
@@ 3.1 Area Graph Definition
The proposed AG is a multi-relational graph [1] ($\mathcal{G}\!=\!\langle\mathcal{V},\mathcal{E}\rangle$) to completely model inter-area relations (Fig. 3). The nodes of AG ($\mathcal{V}$) are image areas, including both areas provided by SAM and areas generated by graph completion (cf. Sec. 3.2). Inspired by classical image pyramid technique, we divide these areas into $L$ levels according to their sizes, corresponding to different image scales, to serve as the foundation of area scale hierarchy. On the other hand, the graph edges ($\mathcal{E}\!=\!\mathcal{E}_{in}\bigcup\mathcal{E}_{adj}$) are used to represent two relations between areas, i.e. inclusion ($\mathcal{E}_{in}$) and adjacency ($\mathcal{E}_{adj}$). The inclusion edge $e_{in}\!\in\!\mathcal{E}_{in}$ is directed, pointing from an area to one of its containing areas. It forms a hierarchical connection between graph nodes, enabling robust and efficient area matching especially under scale changes. In contrast, the adjacency edge $e_{adj}\!\in\!\mathcal{E}_{adj}$ is undirected, indicating the areas it connects share common parts but without the larger one including the smaller one. This edge captures the spatial relations between areas. By the above two edges, AG models both the spatial and scale structure of image areas, thus contributing to exact area matching.
@@ 3.2 Area Graph Construction
The proposed Area Graph takes image areas as nodes, and establishes edges based on area relations. However, not all SAM areas can function as nodes, since some are too small or have extreme aspect ratios, rendering them unsuitable for point matching. Thus, Area Pre-processing is performed first to obtain initial graph nodes. We then approach the edge construction as a Graph Link Prediction problem. Afterwards, the preliminary AG is formed, but it lacks matching efficiency and scale robustness. Thus, we present Graph Completion algorithm, which generates additional nodes and edges to construct the scale hierarchy.
Area Pre-processing. To ensure that SAM areas ($\{a_{i}\big{|}i\in\mathcal{S}\}$) are suitable for area matching, the area pre-processing is adopted first to obtain areas with proper sizes. Specifically, we set two criteria: the acceptable minimal area size ($T_{s}$) and maximum area aspect ratio ($T_{r}$). Any unsuitable area ,$\hat{a}_{i}\big{|}{i}\in\mathcal{S}$, that has smaller size ($W_{\hat{a}_{i}}\times H_{\hat{a}_{i}}<T_{s}$) or larger aspect ratio (${W_{\hat{a}_{i}}}/{H_{\hat{a}_{i}}}>T_{r}$) than thresholds, gets screened out. The remaining areas are added into the candidate areas set $\{a_{i}\big{|}i\in\bar{\mathcal{S}},\bar{\mathcal{S}}\subseteq\mathcal{S}\}$. For each $\hat{a}_{i}$, we fuse it with its nearest neighbor area $a_{n}\big{|}n\in\bar{\mathcal{S}}$, whose center is closer to the center of $\hat{a}_{i}$ than all other areas in the set, to generate a new smallest area $a^{f}$ containing both $\hat{a}_{{i}}$ and $a_{n}$. Then, the $a_{n}$ is replaced by $a^{f}$. We repeat the filtering and fusion on the candidate set until no areas get screened out. Then, we assign a level $l_{a}$ to each candidate area based on its size by setting $L$ size thresholds ($\{TL_{i}~{}\big{|}~{}TL_{i}<TL_{i+1}<2\cdot TL_{i},TL_{0}=T_{s},i\in[0,L-1]\}$):
$l_{a}=i\;\big{|}\;TL_{i}\leq W_{a}\times H_{a}<TL_{i+1}$ (1)
@@
Graph Link Prediction. The edge construction is treated as a link prediction problem. Given two area nodes ($v_{i},v_{j}$), the edge between them ($e_{ij}$) can be predicted according to the spatial relation of their corresponding areas ($a_{i},a_{j}$). This approach adopts the ratio of the overlap size ($O_{ij}$) to the minimum area size between two areas as the score function ($\delta={O_{ij}}/{\min(W_{i}\times H_{i},W_{j}\times H_{j})}$) and can be related to the similarity-based link prediction method [32].
$e_{ij}\in\left\{\begin{aligned} &\mathcal{E}_{in}&,\;&\delta>=\delta_{h}\\ &\mathcal{E}_{adj}&,\;&\delta_{l}<\delta<\delta_{h}\\ &\varnothing&,\;&\delta\leq\delta_{l}\\ \end{aligned}\right.$ (2)
@@
Graph Completion. Initial AG ($\mathcal{G}_{ini}=\langle\mathcal{V}_{ini},\mathcal{E}_{ini}\rangle$) is achieved by connecting all nodes provided by SAM through different edges. However, since SAM inherently produces areas containing complete entity, there are few inclusion relations among areas. Consequently, the $\mathcal{G}_{ini}$ lacks the scale hierarchy, which reduces its robustness at scale variations and makes accessing nodes less efficient. To address the issue, we propose the Graph Completion algorithm, which generates additional nodes and edges to ultimately construct a tree-like structure in the $\mathcal{G}_{ini}$. The core of this algorithm is to generate parent nodes for each orphan node, i.e. the node has no parent node in the next higher level, from small to large size levels. The algorithm begins at the smallest level, collects all orphan nodes on the level, and clusters them based on their center coordinates. Nodes in the same cluster have their corresponding areas fused with their nearest neighbors. It is noteworthy that these generated areas containing multiple objects preserve the internal implicit semantic. Based on our level thresholds, the resulting areas correspond to new higher level nodes, which are added into the $\mathcal{V}_{ini}$ for subsequent operations. If a node remains single after clustering, we increase its area size to the next level to allow for potential parent nodes, in a process known as the upgrade. We repeat the above operations on each level and connect each generated node to others by suitable edges (Eq. 2). More details can be found in the Supp. Mat.
@@ 4 Graphical Area Matching
Since the AG entirely models the global context, the subsequent area matching can be naturally reformulated on the graph. Given two AGs ($\mathcal{G}^{0},\mathcal{G}^{1}$) for the input image pair ($I_{0},I_{1}$) and one area ($a^{0}_{src}\!\in\!I_{0}$) corresponding to node $v^{0}_{src}\!\in\!\mathcal{G}^{0}$ (termed as the source node), area matching involves finding the node $v^{1}_{j}\!\in\!\mathcal{G}^{1}$ with the highest probability of matching its corresponding area $a^{1}_{j}$ to the source node area $a^{0}_{src}$. However, treating this problem as merely independent node matching is inadequate, as it disregards any global information regarding the adjacency and inclusion relations between areas modeled in AG by two kinds of edges, leading to inferior area matches. Meanwhile, these two edges respectively convert AG into two graphical models, i.e. Markov Random Fields (only consider the general adjacency edges) and Bayesian Network (only consider the inclusion edges). Thus, the area matching can be formulated inside a framework of graphical model. Additionally, the complete graphical area matching involves selecting source nodes and exchanging images (cf. LABEL:sec:gamd).
@@ 4.1 Area Markov Random Field
The adjacency edge in AG implies the correlation between probabilities of two nodes matching the source node. By considering the general adjacency relation, which includes the inclusion relations as adjacency too, the $\mathcal{G}^{1}$ is transformed into an undirected graph. The specific conversion involves updating all edges using the following equation:
$e^{1}_{ij}\in\left\{\begin{aligned} &\mathcal{E}_{adj},&\delta>\delta_{l}\\ &\varnothing,&o.w.\\ \end{aligned}\right.$ (3)
Next, random variables ($\bm{x}$) are introduced for all nodes to indicate their matching status with the source node. The binary variable $x_{i}\in\bm{x}$ is equal to $1$ when $v^{1}_{i}$ matches $v^{0}_{src}$ and 00 otherwise. Therefore, Area Markov Random Field (AMRF, $\mathcal{G}^{1}_{M}=\langle\mathcal{V},\mathcal{E}_{adj}\rangle$) is obtained and area matching can be performed by maximizing the joint probability distribution over the AMRF:
$\arg\max_{\bm{x}}P(\bm{x})$ (4)
@@
$E(\bm{x})=\sum_{i}E_{\mathcal{V}}(x_{i})+\lambda\sum_{(i,j)\in\mathcal{N}}E_{\mathcal{E}}(x_{i},x_{j})$ (6)
@@ 4.2 Learning Area Similarity
$Sim_{a^{0},a^{1}}=EX(\sigma_{m}^{0})\times EX(\sigma_{m}^{1})=\sigma_{0,1}\times\sigma_{1,0}$ (9)
Through this approach, the calculation of area similarity is transformed into the patch-level classification, which enhances the interpretability and accuracy of area matching. Its training details can be found in the Supp. Mat.
@@ 4.3 Area Bayesian Network
Although the Graph Cut can be performed in polynomial time [4], the dense energy calculation over $\mathcal{G}_{M}^{1}$ is time consuming. Furthermore, due to the scale hierarchy in AG, this dense calculation is highly redundant. In particular, if the area $a^{0}_{src}$ is not matched to $a^{1}_{j}$, it won’t be matched to any children area of $a^{1}_{j}$. This observation reveals the conditional independence in the similarity calculation:
$\{P(Sim_{a^{0}_{src},a^{1}_{{c}}}=0\big{|}Sim_{a^{0}_{src},a^{1}_{j}}=0)=1\big{|}{c}\in ch^{1}(j)\}$ (10)
where $ch^{i}(j)$ is the index set of children areas of $a^{i}_{j}$ in $\mathcal{G}^{i}$. The conditional independence only involves inclusion edges in $\mathcal{G}^{1}$, thus turning it to a Bayesian Network($\mathcal{G}^{1}_{B}$[3]. Then the redundancy in the similarity calculation can be reduced through $\mathcal{G}^{1}_{B}$. In practice, we calculate the dense similarities by constructing a similarity matrix $M_{S}\!\in\!\mathbb{R}^{|\mathcal{V}^{0}|\times|\mathcal{V}^{1}|}$. Note not all similarities in $M_{S}$ need calculation, but any similarity can be accessed in $M_{S}$. We first calculate similarities related to all source nodes. Subsequent calculations are saved in $M_{S}$ as well. For $M_{S}[i,j]$ that has not been calculated, we obtain it by our learning model (cf. Sec. 4.2).
$M_{S}[i,j]=Sim_{a^{0}_{i},a^{1}_{j}}$ (11)
If $M_{S}[i,j]<T_{as}$, all children nodes $\{v^{0}_{h}\big{|}h\in ch^{0}(i)\}$ and $\{v^{1}_{c}\big{|}c\in ch^{1}(j)\}$ of $v^{0}_{i}$ and $v^{1}_{j}$ are found from $\mathcal{G}_{B}^{0}$ and $\mathcal{G}_{B}^{1}$. Based on the conditional independence, we can get:
$\{M_{S}[h,k]=0\big{|}h\in ch^{0}(i),k\in ch^{1}(j)\}$ (12)

### 08 Apr 2024 v2

@@ 1 Introduction
The underlying goal of feature matching is to find correspondences for points in the source image from a highly redundant candidate set, i.e. all the points of the target image. Therefore, feature matching essentially involves reducing the matching redundancy, which is caused by troublesome matching noises. Recent years have witnessed significant advancements in learning-based methods for feature matching. Especially the semi-dense [40, 8] and dense methods [17], which search for matches densely by deep feature comparison, obtain an impressive precision gap over sparse methods [14]. However, as they heavily rely on dense feature comparison of the entire image, limited input resolution and error-prone computation in matching-redundant parts of images decrease their accuracy. Thus, effective matching redundancy reduction is urgently required for them.
On the other hand, most matching redundancy can be effectively and efficiently identified through high-level image understanding and only strongly correlated local areas (also known as regions) need dense feature comparison to determine precise matches. (Fig. 1). Therefore, recent methods [15, 9] perform learning-based overlap segmentation before matching. However, implicit learning needs non-reusable computational overhead and the redundancy still exists in the overlap. To address these issues, some work turn to explicit semantic prior [18, 52]. Unlike manually specifying a topic number to group feature patches [18], Area to Point Matching (A2PM) framework proposed by SGAM [52] provides a more intuitive way to reduce matching redundancy. Specifically, this framework establishes semantic area matches first where matching redundancy is largely removed, and then obtains accurate correspondences within these areas utilizing an state-of-the-art point matcher. Nevertheless, SGAM relies on semantic segmentation to determine area matches. Its performance, thus, decreases when encountering inexact semantic labeling and semantic ambiguity [52]. Furthermore, the benefits of A2PM cannot be extrapolated to more general scenes due to the close-set semantic labels. Hence, reducing matching redundancy through semantic segmentation suffers from impracticality.
@@
In this work, we propose Matching Everything by Segmenting Anything (MESA, Fig. 2), a method for precise area matching from SAM segmentation. MESA focuses on two main aspects: constructing area relations and finding area matches based on these relations. Since SAM areas only provide local information, matching them independently can lead to inaccurate results, especially in scenes with scale variation and repetitiveness. To address this, MESA first constructs a novel graph structure, named Area Graph (AG), to model the global context of areas for subsequent precise matching. The AG takes areas as nodes and connects them with two types of edges: undirected edges for adjacent areas and directed edges for areas that include one another. Both edges capture global information, and the latter also enables the construction of hierarchy structures in AG similar to [43] for efficient matching. Afterwards, MESA performs area matching by deriving two distinct graphical models from the AG: Area Markov Random Field (AMRF) and Area Bayesian Network (ABN). The AMRF involves global-informative edges in AG, thus allowing global-consistent area matching through energy minimization on the graph. The ABN, furthermore, is proposed to facilitate the graph energy calculation, through the hierarchy structure of AG. Specifically, the graph energy is determined based on the similarity and spatial relation between areas, making this energy minimization effectively solvable through the Graph Cut [23]. The area similarity calculation is accomplished by a novel learning-based model and accelerated by the ABN. Finally, we propose a global matching energy to refine accurate area matches, addressing the issue of multiple solutions in Graph Cut, leading to effective matching redundancy reduction.
@@
Figure 2: Overview of MESA. Based on ❶ SAM segmentation, we first construct ❷ Area Graphs. Then the graph is turned to two graphical models based on its two different edges. Through ❸ Area Markov Random Field, area matching is formulated as an ❹ Energy Minimization. Then, leveraging ❺ Area Bayesian Network and our ❺ Learning Area Similarity Calculation, ❻ Graph Energy can be efficiently calculated. Therefore, ❼ Graph Cut is utilized to obtain putative area matches. Finally, ❽ Global Energy Minimization determines the best area match, which serves as the input of subsequent point matcher for precise feature matching, following the ❾ Area to Point Matching framework [52].
@@ 3 Area Graph
In this section, we introduce the concept of an Area Graph (AG) and explain how to construct the AG from the SAM [22] results of an image. The main reason to adopt the AG is that direct area matching on SAM results is unacceptable, as global information is ignored in independent areas. Also, fixed area sizes hinder robust point matching under scale changes. Therefore, AG is proposed to capture the global structural information of these areas and construct scale hierarchy for them, contributing to accurate and robust area matching.
@@ 3.1 Area Graph Definition
The AG ($\mathcal{G}\!=\!\langle\mathcal{V},\mathcal{E}\rangle$) takes image areas as nodes and contains two kinds of edges to model inter-area relations (Fig. 3), thus making it a multi-relational graph [1]. The graph nodes include both areas provided by SAM and areas generated by graph completion (cf. Sec. 3.2). We divide these areas into $L$ levels according to their sizes, corresponding to different image scales, to serve as the foundation of area scale hierarchy. On the other hand, the graph edges ($\mathcal{E}\!=\!\mathcal{E}_{in}\bigcup\mathcal{E}_{adj}$) represent two relations between areas, i.e. inclusion ($\mathcal{E}_{in}$) and adjacency ($\mathcal{E}_{adj}$). The inclusion edge $e_{in}\!\in\!\mathcal{E}_{in}$ is directed, pointing from an area to one of its containing areas. It forms a hierarchical connection between graph nodes, enabling robust and efficient area matching especially under scale changes. The adjacency edge $e_{adj}\!\in\!\mathcal{E}_{adj}$ is undirected, indicating the areas it connects share common parts but without the larger one including the smaller one. This edge captures the spatial relations between areas. By the above two edges, AG models both the spatial and scale structure of image areas, thus contributing to exact area matching.
@@ 3.2 Area Graph Construction
The construction of AG includes collecting areas as nodes and connecting them by proper edges. Firstly, not all SAM areas can function as nodes, since some are too small or have extreme aspect ratios, rendering them unsuitable for point matching. Thus, Area Pre-processing is performed first to obtain initial graph nodes. We then approach the edge construction as a Graph Link Prediction problem [34]. Afterwards, the preliminary AG is formed, but it still lacks matching efficiency and scale robustness. Thus, we propose the Graph Completion algorithm, which generates additional nodes and edges to construct the scale hierarchy.
Area Pre-processing. To filter unsuitable areas, we set two criteria: the acceptable minimal area size ($T_{s}$) and maximum area aspect ratio ($T_{r}$). Any area that has smaller size than $T_{s}$ or larger aspect ratio than $T_{r}$, gets screened out. The remaining areas are added into the candidate areas set. For each filtered area, we fuse it with its nearest neighbor area in the candidate set. We repeat the filtering and fusion on the candidate set until no areas get screened out. Then, we assign a level $l_{a}$ to each candidate area $a$ based on its size, by setting $L$ size thresholds ($\{TL_{i}~{}\big{|}~{}i\in[0,L-1]\}$):
$l_{a}=i\;\big{|}\;TL_{i}\leq W_{a}\times H_{a}<TL_{i+1}.$ (1)
@@
Graph Link Prediction. The edge construction is treated as a link prediction problem [34]. Given two area nodes ($v_{i},v_{j}$), the edge between them ($e_{ij}$) can be predicted according to the spatial relation of their corresponding areas ($a_{i},a_{j}$). This approach adopts the ratio of the overlap size ($O_{ij}$) to the minimum area size between two areas ($\delta={O_{ij}}/{\min(W_{i}\times H_{i},W_{j}\times H_{j})}$) as the score function:
$e_{ij}\in\left\{\begin{aligned} &\mathcal{E}_{in}&,\;&\delta>=\delta_{h}\\ &\mathcal{E}_{adj}&,\;&\delta_{l}<\delta<\delta_{h}\\ &\varnothing&,\;&\delta\leq\delta_{l}\\ \end{aligned}\right.,$ (2)
@@
Graph Completion. Initial AG is achieved by connecting all processed nodes with different edges. However, since SAM inherently produces areas containing complete entity, there are few inclusion relations among areas. Consequently, initial AG lacks the scale hierarchy, which reduces its robustness at scale variations and makes accessing nodes inefficient. To address the issue, we propose the Graph Completion algorithm, which generates additional nodes and edges to ultimately construct a tree structure in the original graph. The core of this algorithm is to generate parent nodes for each orphan node, i.e. the node has no parent node in the next higher level, from small to large size levels. The algorithm begins at the smallest level, collects all orphan nodes, and clusters them based on their center coordinates. Nodes in the same cluster have their corresponding areas fused with their nearest neighbors. It is noteworthy that these generated areas containing multiple objects preserve the internal implicit semantic. Based on our level thresholds, the resulting areas correspond to new higher level nodes. If a node remains single after clustering, we increase its area size to the next level to allow for potential parent nodes. We repeat the above operations on each level and connect generated nodes to others by suitable edges. More details can be found in the Supp. Mat.
@@ 4 Graphical Area Matching
In this section, we describe our area matching method, which formulates area matching on the graph, utilizing two graphical models derived from the AG. Given two AGs ($\mathcal{G}^{0},\mathcal{G}^{1}$) of the input image pair ($I_{0},I_{1}$) and one area ($a^{0}_{src}\!\in\!I_{0}$) corresponding to node $v^{0}_{src}\!\in\!\mathcal{G}^{0}$ (termed as the source node), area matching involves finding the node $v^{1}_{j}\!\in\!\mathcal{G}^{1}$ with the highest probability of matching its corresponding area $a^{1}_{j}$ to the source node area $a^{0}_{src}$. However, treating this problem as independent node matching is inadequate, as which disregards global information from adjacency and inclusion relations between areas modeled in AG by its two edges. Meanwhile, these two edges respectively convert AG into two graphical models, i.e. Markov Random Fields (undirected edges) and Bayesian Network (directed edges). Thus, the area matching can be naturally formulated inside a framework of graphical model.
@@ 4.1 Area Markov Random Field
The adjacency edge in AG implies the correlation between probabilities of two nodes matching the source node. By considering the general adjacency relation, which includes the inclusion relations as adjacency too, the $\mathcal{G}^{1}$ is transformed into an undirected graph. Next, random variables ($\bm{x}$) are introduced for all nodes to indicate their matching status with the source node. The binary variable $x_{i}\in\bm{x}$ is equal to $1$ when $v^{1}_{i}$ matches $v^{0}_{src}$ and 00 otherwise. Therefore, Area Markov Random Field (AMRF, $\mathcal{G}^{1}_{M}=\langle\mathcal{V},\mathcal{E}_{adj}\rangle$) is obtained and area matching can be performed by maximizing the joint probability distribution over the AMRF:
$\arg\max_{\bm{x}}P(\bm{x}).$ (3)
@@
$E(\bm{x})=\sum_{i}E_{\mathcal{V}}(x_{i})+\lambda\sum_{(i,j)\in\mathcal{N}}E_{\mathcal{E}}(x_{i},x_{j}),$ (5)
@@ 4.2 Learning Area Similarity
$Sim_{a^{0},a^{1}}=EX(\sigma_{m}^{0})\times EX(\sigma_{m}^{1})=\sigma_{0,1}\times\sigma_{1,0}.$ (8)
The training details of this learning model are described in the Supp. Mat.
@@ 4.3 Area Bayesian Network
Although the Graph Cut can be performed in polynomial time [4], the dense energy calculation over $\mathcal{G}_{M}^{1}$ is time consuming. Furthermore, due to the scale hierarchy in AG, this dense calculation is highly redundant. In particular, if the area $a^{0}_{src}$ is not matched to $a^{1}_{j}$, it won’t be matched to any children area of $a^{1}_{j}$. This observation reveals the conditional independence in the similarity calculation, which involves inclusion edges in $\mathcal{G}^{1}$, thus turning $\mathcal{G}^{1}$ to a Bayesian Network ($\mathcal{G}^{1}_{B}$[3]. Therefore, the redundancy in the similarity calculation can be reduced. In practice, we calculate the dense similarities by constructing a similarity matrix $M_{S}\!\in\!\mathbb{R}^{|\mathcal{V}^{0}|\times|\mathcal{V}^{1}|}$. Note not all similarities in $M_{S}$ need calculation, but any similarity can be accessed in $M_{S}$. We first calculate similarities related to all source nodes. Subsequent calculations are saved in $M_{S}$ as well. For $M_{S}[i,j]$ that has not been calculated, we achieve it by our learning model (cf. Sec. 4.2):
$M_{S}[i,j]=Sim_{a^{0}_{i},a^{1}_{j}}.$ (9)
If $M_{S}[i,j]<T_{as}$, all children nodes $\{v^{0}_{h}\big{|}h\in ch^{0}(i)\}$ and $\{v^{1}_{c}\big{|}c\in ch^{1}(j)\}$ of $v^{0}_{i}$ and $v^{1}_{j}$ are found from $\mathcal{G}_{B}^{0}$ and $\mathcal{G}_{B}^{1}$, where $ch^{j}(i)$ is the index set of children indices of node $v_{i}^{j}$ from $\mathcal{G}^{j}_{B}$. Based on the conditional independence, we have:
$M_{S}[h,k]=0,~{}\forall(h,k)\in ch^{0}(i)\times ch^{1}(j).$ (10)